Need a little help. Please help

You wish to test the claim that the average IQ score is less than 100 at the .10 significance level. You determine the hypotheses are:

Ho : μ = 100 H 1 : μ < 100

You take a simple random sample of 41 individuals and find the mean IQ score is 97.2, with a standard deviation of 14.9. Let's consider testing this hypothesis two ways: once with assuming the population standard deviation is not known and once with assuming that it is known.

Round to three decimal places where appropriate.

Assume Population Standard Deviation is NOT known Assume Population Standard Deviation is 15
Test Statistic: t = Test Statistic: z =
Critical Value: t = Critical Value: z =
p-value: p-value:
Conclusion About the Null: (pick one)
  • Reject the null hypothesis
  • Fail to reject the null hypothesis
Conclusion About the Null: (pick one)
  • Reject the null hypothesis
  • Fail to reject the null hypothesis
Conclusion About the Claim: (pick one)
  • There is sufficient evidence to support the claim that the average IQ score is less than 100.
  • There is NOT sufficient evidence to support the claim that the average IQ score is less than 100.
  • There is sufficient evidence to warrant rejection of the claim that the average IQ score is less than 100.
  • There is NOT sufficient evidence to warrant rejection of the claim that the average IQ score is less than 100.
Conclusion About the Claim: (pick one)
  • There is sufficient evidence to support the claim that the average IQ score is less than 100.
  • There is NOT sufficient evidence to support the claim that the average IQ score is less than 100.
  • There is sufficient evidence to warrant rejection of the claim that the average IQ score is less than 100.
  • There is NOT sufficient evidence to warrant rejection of the claim that the average IQ score is less than 100.

Is there a significant difference between when we know the population standard deviation and when we don't? Explain.
Answer #1
Assume Population Standard Deviation is NOT known

Sample size = n = 41

Sample mean = T = 97.2

Standard deviation = s = 14.9

Here population standard deviation is unknown so we have to use t-test statistic.
Test statistic is

t Vn.

97.2 100 -2.8 -1.20 14.9 2.326989 V41

Degrees of freedom = n - 1 = 41 - 1 = 40

Level of signficance = 0.10

Critical value = 1.303

P-value = P(T < - 1.20) = 0.1180

P-value > 0.10 we fail to reject null hypothesis.

  • There is NOT sufficient evidence to support the claim that the average IQ score is less than 100.

------------------------------------------------------------------------------------------------------------------------------------------------------

Assume Population Standard Deviation is 15

Sample size = n = 41

Sample mean = T = 97.2

Population standard deviation = \sigma = 15

Here population standard deviation is known so we have to use z-test statistic.

Test statistic is

Τμ υ σ//n

97.2 100 -2.8 -1.20 15/V41 2.342606

Critical value = 1.28

P-value = P(Z < - 1.20) = 0.1160

P-value > 0.10 we fail to reject null hypothesis.

  • There is NOT sufficient evidence to support the claim that the average IQ score is less than 100.